Determine whether (-2,3)is a solution of the equation x+3y=11
Found 2 solutions by jim_thompson5910, harpazo:
Answer by jim_thompson5910(35256)

The given point is (-2,3). So, (x,y) = (-2,3) which means x = -2 and y = 3

Plug in the given x and y values. If we get a true equation, then the ordered pair is a solution. If not, then we don't have a solution.

x+3y=11

-2+3*(3)=11 ... plug in x = -2 and y = 3

-2+9=11

7 = 11 ... this equation is FALSE

Since the last equation is FALSE, this means that (-2,3) is NOT a solution to the equation x+3y=11

Answer by harpazo(655)

Plug -2 for x and 3 for y in the given equation.
If after doing the math, you get the same answer on both sides of the equation, you will then know that the point given is a solution.
In other words, do the math and try to get 11 = 11.
You do the math!